The pH of a 0.0516 M solution of nitrous acid, HNO2, is 2.34. What is its Ka?
1
Expert's answer
2019-07-01T03:54:46-0400

Solution.

Nitrous acid is a very weak acid, so you need to write the following formula:


"pH = \frac{pKa}{2} - \frac{lg(H+)}{2}"

"pKa = 2 \times (pH + \frac{lg(H+)}{2})"

"pKa = 2 \times (2.34 + \frac{lg(0.0516)}{2})"

pKa = 3.39


"pKa = -lg(Ka)"

"Ka = 10^{-pKa}"

"Ka = 4.07 \times 10^{-4}"

Answer:

"Ka = 4.07 \times 10^{-4}"